Prove that the points A(2,-2) , B(8,4) , C(5,7) , D(-1,1) are the vertices of a rectangle
 

Grade:11

1 Answers

Arun
4 years ago
 
Dear student
 
As AB =sqrt (72) = CD
 
BC = sqrt (18) = DA
 
Also AC = BD = sqrt (90)
 
hence ABCD will be rectangle.